OpenStudy (anonymous):

The half life of a certain isotope is 2 yrs a sample of 3200 kgs of the isotopes is considered if x kgs of sample remains after 12 yrs then x =?

OpenStudy (anonymous):

dude ur funny and creepy

OpenStudy (anonymous):

Why???

OpenStudy (anonymous):

12 years is \(12\div2=6\) half lives

OpenStudy (unklerhaukus):

\[A(t)=A_02^{-t/t_{1/2}}\]

OpenStudy (anonymous):

solve via \(3200\times \frac{1}{2^6}\)

OpenStudy (unklerhaukus):

\[A(t)=A_02^{-t/t_{1/2}}\] \[\downarrow\] \[X(12)=3200\cdot2^{-12/2}\]

OpenStudy (anonymous):

Can u guys tell me wat u r soing

OpenStudy (anonymous):

where that went......

OpenStudy (anonymous):

my arithmetic is bad let me try again start with \(3200\) and the half life is 2 years so in two years there will be \[3200\times \frac{1}{2}=1600\]

OpenStudy (anonymous):

in another two years there will be half of that, so \[1600\times \frac{1}{2}=800\] in another two years half of that, so \[800\times \frac{1}{2}=400\] and in another two years half that so \[400\times\frac{1}{2}=200\] and so on multiply by \(\frac{1}{2}\) for each half life

OpenStudy (anonymous):

it is much easier to do this in one step 12 years is 6 half lives, so you can compute in one step via \[3200\times \left(\frac{1}{2}\right)^6\]

OpenStudy (ujjwal):

After n half lives, the fraction of radioactive element left is:\[(\frac{1}{2})^n\]

OpenStudy (anonymous):

rather than dividing by two repeatedly

OpenStudy (anonymous):

2^6 = 64

OpenStudy (anonymous):

=50

OpenStudy (anonymous):

Yup......thxxx

OpenStudy (unklerhaukus):

\[X=3200\cdot2^{-12/2}\]\[=2^5\times10^2\times2^{-6}\]\[=2^{-1}\times10^2\]\[=\frac12\times100\]